# Rec-RIR: Monaural Blind Room Impulse Response Identification via DNN-based Reverberant Speech Reconstruction in STFT Domain

### v1

@@ 4.1 Datasets
We use the SimACE test set proposed in [12], in which the microphone recordings are generated by convolving the clean speech from the ’si_et_05’ subset in WSJ0 corpus [26] with the downsampled measured RIRs from the ’Single’ subset in ACE Challenge [27], and adding noise from the test set in REVERB Challenge with a SNR of 20 dB. The reference minimum and maximum RT60s are 0.332 s and 1.22 s, respectively [27]. During testing, the labels of acoustic parameters are computed from the ground truth RIRs [12].
4.2 Settings
Before Rec-RIR, the input recordings are normalized using their maximum absolute value. The STFT analysis and synthesis windows are square-root Hann windows with a length of 512 samples and 50% overlap, which means $F=257$. The numbers of layers in the proposed DNN are $M_{\mathrm{noi}}=2$, $M_{\mathrm{rev}}=6$ and $M_{\mathrm{ctf}}=4$. The dimension of embeddings is set to $C=96$. The CTF length is set to $L=60$, which corresponds to RIR estimates with an effective length of approximately 0.96 s. With the above settings, Rec-RIR has 3.1 M parameters and 62.2 GFlops/s computational complexity.
For training, we set $\lambda_{\mathrm{rvb}}=1$ and $\lambda_{\mathrm{cln}}=1$, segment speech utterances into 4 s, and use 97,092 samples per epoch with a batch size of 4. We adopt the AdamW optimizer [28] and the cosine annealing learning rate scheduler with restarts. The learning rate starts at 0.001 per epoch, decays cosinely, and reaches its minimum at the end of epoch. Training runs for 35 epochs total.
@@ 4.2 Settings
4.3 Comparison methods and metrics
@@
We evaluate Rec-RIR using acoustic parameters RT60, DRR and C50. We present their mean absolute error (MAE), root mean square error (RMSE), and Pearson correlation coefficient ($\rho$). Additionally, we provide the RMSE of early reflections (marked as RIR-50 ms), corresponding to the 50 ms duration after the direct-path impulse.
4.4 Results and analysis
As shown in Table 1, Rec-RIR achieves the SOTA estimation of the early reflections and the acoustic parameters. The Pearson correlation coefficient of Rec-RIR is close to 1, indicating a more stable and consistent estimation compared to other methods. Unlike FiNS which learns the early reflections and the noise shaping filters directly from fixed-length inputs, BUDDy, VINP, and Rec-RIR solve the filter in STFT domain based on the spectra of reverberant recordings with arbitrary length. Among these three approaches, the DNNs in BUDDy and VINP do not directly output the filter coefficients. Instead, they alternately update the estimates of clean speech and filters according to the signal model and derive the final estimates after multiple iterations by maximizing likelihood. Differently, Rec-RIR introduces the DNN and loss functions specifically designed for RIR estimation, enabling a direct estimation of CTF filter without the need for any iteration, thus reducing the computational complexity. Fig. 3 shows an example of the estimated RIR, whose ground truth RT60, DRR and C50 are 1.26 s, -1.24 dB and 4.90 dB, respectively. It can be found that Rec-RIR is capable of estimating the impulses in early reflections. The impulses in the measured RIR that occur before 2 ms cannot be fully reconstructed, which is reasonable because these impulses correspond to the frequency response of the measurement system. During training, we use the direct-path signal as the clean speech to obtain the aligned time delay and gain. However, since the direct-path signal contains the frequency response, the estimated CTF filter may lack the corresponding component.
Fig. 3: Example of ground truth and estimated RIRs.
@@ 4.4 Results and analysis
Fig. 4: 2-D projections of CTF embeddings on SimACE.
4.5 Ablation study
An ablation study with variant loss weights is designed to demonstrate the contribution of the auxiliary losses. We present the mean values and standard deviations of metrics in Table 2. Compared with Var-1 which estimates the CTF filter directly using the primary loss, other experiments have lower errors, indicating that both auxiliary losses contribute to the final performance. The MAEs of RT60 for Var-2, Var-3 and Rec-RIR are close. We further employ the two-tailed Z-test to analyze their differences. Results show that there are no significant statistical differences in RT60 estimation between Rec-RIR and Var-2 as well as Var-3 with p-values $p>0.999$ and $p\approx 0.306>0.05$, respectively. Therefore, the proposed setting is the best among all evaluated settings.
Table 2: Ablation study for loss functions.
@@ 4.5 Ablation study
5 conclusion
In this work, we propose Rec-RIR for blind RIR identification. According to CTF approximation, we propose an end-to-end DNN that consists of cross-band and narrow-band blocks with auxiliary losses to estimate the CTF filter by reconstructing the reverberant speech spectrum in the STFT domain. A pseudo intrusive measurement process is employed to simulate the common intrusive RIR measurement process and transform the CTF filter into RIR. Experimental results show that Rec-RIR attains SOTA estimation of RIR and acoustic parameters.

### v2

@@ IV-A Datasets
We use the SimACE test set proposed in [29], in which the microphone recordings are generated by convolving the clean speech from WSJ0 corpus [14] with the downsampled measured RIRs from ACE Challenge [5], and adding noise from the test set in REVERB Challenge with a SNR of 20 dB. During testing, the labels of the acoustic parameters are calculated from the ground truth RIRs.
IV-B Settings
Before Rec-RIR, the input recordings are normalized using their maximum absolute value. The STFT analysis and synthesis windows are square-root Hann windows with a length of 512 samples and 50% overlap, which means $F=257$. The numbers of layers are set to $M_{\mathrm{noi}}=2$, $M_{\mathrm{rev}}=6$ and $M_{\mathrm{ctf}}=6$. The dimension of embeddings is set to $C=96$. The CTF length is set to $L=60$, which corresponds to RIR estimates with an effective length of approximately 0.96 s. With the above settings, Rec-RIR has 3.1 M parameters and is with 35.2 GMACs/s computational complexity.
For training, we set $\lambda_{\mathrm{rvb}}=1$ and $\lambda_{\mathrm{cln}}=1$, segment speech utterances into 4 s, and use 97,092 samples per epoch with a batch size of 4. We adopt the AdamW optimizer [12] and the cosine annealing learning rate scheduler with restarts. The learning rate starts at 0.001 and decays cosinely in each epoch. Training runs for 35 epochs.
@@ IV-B Settings
IV-C Comparison methods and metrics
@@
We evaluate Rec-RIR using acoustic parameters RT60, DRR and C50. We present their mean absolute error (MAE), root mean square error (RMSE), and Pearson correlation coefficient ($\rho$). Additionally, we provide RMSE and average $\rho$ ($\bar{\rho}$) of early reflections (marked as RIR-50 ms), corresponding to the 50 ms duration after the direct-path impulse.
IV-D Results and analysis
As shown in Table I, Rec-RIR achieves the SOTA estimation of the early reflections and the acoustic parameters. The high Pearson correlation coefficients of Rec-RIR indicate stable RIR identification results. Unlike FiNS which learns the early reflections and the noise shaping filters directly from fixed-length inputs, BUDDy, VINP, and Rec-RIR solve the filter in STFT domain based on the spectra of reverberant recordings with arbitrary length. Among these three approaches, the DNNs in BUDDy and VINP do not directly output the filter coefficients. Instead, they alternately update the estimates of clean speech and filters according to the signal model and derive the final estimates after multiple iterations by maximizing likelihood. Differently, Rec-RIR introduces the DNN and loss functions specifically designed for RIR estimation, enabling a direct estimation of CTF filter without the need for any iteration, thus reducing the computational complexity. Fig. 3 shows an example of the estimated RIR, whose ground truth RT60, DRR and C50 are 1.26 s, -1.24 dB and 4.90 dB, respectively. It can be found that Rec-RIR is capable of estimating the impulses in early reflections. The impulses in the measured RIR that occur before 2 ms cannot be fully reconstructed, which is reasonable because we use the direct-path signal as the clean speech. The direct-path signal constrains the time delay and gain of the RIR, resolving the underdetermined problem in RIR identification. However, during training, since the direct-path signal inevitably contains part of frequency response, the estimated RIR may lack the corresponding component.
Figure 3: Example of ground truth and estimated RIRs.
@@ IV-D Results and analysis
Figure 4: 2-D projections of CTF embeddings on SimACE.
IV-E Ablation study
An ablation study with variant loss weights is designed to demonstrate the contribution of the auxiliary losses. We present the mean values and standard deviations of metrics in Table II. Compared with Var-1 which estimates the CTF filter directly using the primary loss, other experiments have lower errors, indicating that both auxiliary losses contribute to the final performance. The MAEs of RT60 for Var-2, Var-3 and Rec-RIR are close. We further employ the two-tailed Z-test to analyze their differences. Results show that there are no significant statistical differences in RT60 estimation between Rec-RIR and Var-2 as well as Var-3 with p-values $p>0.999$ and $p\approx 0.306>0.05$, respectively. Therefore, the proposed setting is the best among all evaluated settings.
TABLE II: Ablation study for loss functions.
@@ IV-E Ablation study
V conclusion
We propose Rec-RIR for blind RIR identification. In Rec-RIR, a DNN that consists of cross-band and narrow-band blocks with auxiliary losses is proposed. Rec-RIR estimates the CTF filter directly by reconstructing the noise-free reverberant speech spectra in the STFT domain. After that, a pseudo intrusive measurement process is employed to convert the CTF filter into RIR. Experimental results show that Rec-RIR attains SOTA estimation of RIR and acoustic parameters.